OpenStudy (anonymous):

Find the intercepts: y=(2x-1)/(3-x)

jimthompson5910 (jim_thompson5910):

To find the y intercept (there is only one y-intercept), plug in x = 0 and evaluate y=(2x-1)/(3-x) y=(2(0)-1)/(3-0) y=(-1)/(3) y = -1/3 So the y-intercept is (0, -1/3)

OpenStudy (anonymous):

Plug in x =0 to find y intercept y = 2x -1 / 3- x y = 0-1/ 3-0 = -1/3 Y intercept is (0,-1/3) Plug in y = 0 and find x y=(2x-1)/(3-x) 0 = (2x-1)/(3-x) Solve for x 2x -1 =0 x = 1/2 X intercept is (1/2 , 0)

jimthompson5910 (jim_thompson5910):

To find the x -intercept, plug in y = 0 and solve for x y=(2x-1)/(3-x) 0=(2x-1)/(3-x) 2x-1 = 0 2x = 1 x = 1/2 So the x-intercept is (1/2, 0)